# Expressive Power of Graph Neural Networks for (Mixed-Integer) Quadratic Programs

### v1

@@ 4.1. GNNs cannot universally represent MI-LCQPs
In this subsection, we present some counter-examples illustrating the fundamental limitation of GNNs for representing properties of MI-LCQPs. It is known in the previous literature that there exists some continuous function that cannot be approximated by GNNs with arbitrarily small error, see e.g., [xu2019powerful, azizian2020expressive, geerts2022expressiveness], and our results in Section 3 indicates that properties of LCQPs do not suffer from this fundamental limitation. However, when integer variables are introduced, GNN will fail to distinguish some pairs of MI-LCQP-graphs with different properties. The first pair of counter-examples (feasible vs infeasible) is directly from [chen2022representing-milp] for MILPs which also allies for MI-LCQPs since adding a quadratic term in the objective function of an MILP problem does not change the feasible region.
Proposition 4.1 ([chen2022representing-milp]*Lemma 3.2).
@@
The next two propositions, not covered in [chen2022representing-milp], show that GNNs have limited separation power to distinguish two feasible MILP problems with different optimal objective values and two feasible MILP problems with the same optimal objective values and disjoint optimal solution sets.
@@
Propositions 4.1, 4.2, and 4.3 indicate that there exist some MI-LCQP data distribution on which it is impossible to train a GNN to predict MI-LCQP properties, regardless of the size or the complexity of the GNN. In particular, one can just choose the uniform distribution over pairs of instances satisfying Propositions 4.1, 4.2, and 4.3, for which any GNN making good approximation on one instance must fail on the other.
We explicitly present a pair of MI-LCQP instances that prove Proposition 4.2 in Figure 2 while deferring the examples for Propositition 4.3 to Appendix B.
$\begin{split}\min_{x\in\mathbb{R}^{6}}&~{}~{}\frac{1}{2}\sum_{i=1}^{6}x_{i}^{2}+\sum_{i=1}^{6}x_{i},\\ \text{s.t.}&~{}~{}x_{1}+x_{2}\geq 1,\ x_{2}+x_{3}\geq 1,\ x_{3}+x_{4}\geq 1,\\ &~{}~{}x_{4}+x_{5}\geq 1,\ x_{5}+x_{6}\geq 1,\ x_{6}+x_{1}\geq 1,\\ &~{}~{}x_{j}\in\{0,1\},\ \forall~{}j\in\{1,2,\dots,6\}.\end{split}$
$\begin{split}\min_{x\in\mathbb{R}^{6}}&~{}~{}\frac{1}{2}\sum_{i=1}^{6}x_{i}^{2}+\sum_{i=1}^{6}x_{i},\\ \text{s.t.}&~{}~{}x_{1}+x_{2}\geq 1,\ x_{2}+x_{3}\geq 1,\ x_{3}+x_{1}\geq 1,\\ &~{}~{}x_{4}+x_{5}\geq 1,\ x_{5}+x_{6}\geq 1,\ x_{6}+x_{4}\leq 1\\ &~{}~{}x_{j}\in\{0,1\},\ \forall~{}j\in\{1,2,\dots,6\}.\\ ~{}\end{split}$
$\min~{}\frac{1}{2}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}&{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}&{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}\begin{bmatrix}1&0.1&\\ 0.1&1&\\ &&1\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}+\begin{bmatrix}1&2&3\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}$s.t.
$2{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}$ $+$ ${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $\leq 5$
${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $+$ $3{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}$ $\geq 0$
$0\leq{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}},{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}},{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\leq 1$
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$
$\min~{}\frac{1}{2}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}&{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}&{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}\begin{bmatrix}1&0.1&\\ 0.1&1&\\ &&1\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}+\begin{bmatrix}1&2&3\end{bmatrix}\begin{bmatrix}{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}\\ {\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}\\ {\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\end{bmatrix}$s.t.
$2{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}}$ $+$ ${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $\leq 5$
${\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}}$ $+$ $3{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}$ $\geq 0$
$0\leq{\color[rgb]{0.203125,0.59765625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.203125,0.59765625,1}x_{1}},{\color[rgb]{1,0.296875,0.234375}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.296875,0.234375}x_{2}},{\color[rgb]{0.1953125,0.78515625,0.1953125}\definecolor[named]{pgfstrokecolor}{rgb}{0.1953125,0.78515625,0.1953125}x_{3}}\leq 1$
$v_{1}$$v_{2}$$v_{3}$$v_{4}$$v_{5}$$v_{6}$$w_{1}$$w_{2}$$w_{3}$$w_{4}$$w_{5}$$w_{6}$
Figure 2. MI-LCQP instances proving Proposition 4.2
We can observe that the two MI-LCQP-graphs are not isomorphic, but all vertices of the same color have the same information from neighbors. For example, each red vertex $w_{j}$ has two blue neighbors with $A_{ij}=1$ and is connected to itself with a self-loop $Q_{jj}=1$. Therefore, even if they are not isomorphic, they lead to the same output for any GNN in $\mathcal{F}_{\textup{MI-LCQP}}$, for which we include a detailed proof in Appendix B. However, they have different properties. Particularly, they are both feasible but have different optimal objective values. More specifically, the optimal objective value of the first instance is $\frac{9}{2}$ that can be achieved at $(1,0,1,0,1,0)$. For the second instance, the constraints $x_{1}+x_{2}\geq 1$, $x_{2}+x_{3}\geq 1$, $x_{3}+x_{1}\geq 1$, and $x_{1},x_{2},x_{3}\in\{0,1\}$ imply that at least two of $x_{1},x_{2},x_{3}$ take the value of $1$. A similar analysis also applies to $x_{4},x_{5},x_{6}$, and hence, the optimal objective value of the second instance is $6\neq\frac{9}{2}$.
@@ 4.2. GNNs can represent particular types of MI-LCQPs
We have shown a fundamental limitation of GNNs to represent properties of general MI-LCQP problems. Therefore, a natural question is whether we can identify a subset of $\mathcal{G}_{\textup{MI-LCQP}}$ on which it is possible to train reliable GNNs. To address this, we need to gain a better understanding for the separation power of GNNs or equivalently the WL test, according to the discussion following Theorem 3.4. We state in Algorithm 1 the WL test for MI-LCQP-graphs associated to $\mathcal{F}_{\textup{MI-LCQP}}$ or $\mathcal{F}_{\textup{MI-LCQP}}^{W}$, where $C_{i}^{l,V}$ and $C_{j}^{l,W}$ are understood as the color of $i\in V$ and $j\in W$ at the $l$-th iteration.
Algorithm 1 The WL test for MI-LCQP-graphs
1:A LCQP-graph $G=(V,W,A,Q,H_{V},H_{W})$ and iteration limit $L>0$.
@@
4:     $C_{i}^{l,V}=\text{HASH}\big{(}C_{i}^{l-1,V},\big{\{}\big{\{}(C_{j}^{l-1,W},A_{ij}):j\in\mathcal{N}_{i}^{W}\big{\}}\big{\}}\big{)}$.
5:     $C_{j}^{l,W}=\text{HASH}\big{(}C_{j}^{l-1,W},\big{\{}\big{\{}(C_{i}^{l-1,V},A_{ij}):i\in\mathcal{N}_{j}^{V}\big{\}}\big{\}},\big{\{}\big{\{}(C_{j^{\prime}}^{l-1,W},Q_{jj^{\prime}}):j^{\prime}\in\mathcal{N}_{j}^{W}\big{\}}\big{\}}\big{)}.$
@@
7:return The multisets containing all colors $\big{\{}\big{\{}C_{i}^{L,V}\big{\}}\big{\}}_{i=0}^{m},\big{\{}\big{\{}C_{j}^{L,W}\big{\}}\big{\}}_{j=0}^{n}$.
@@
Initially, each vertex is labeled a color according to its attributes ($v_{i}$ or $w_{j}$). In the case that the hash functions introduce no collisions, two vertices are of the same color at the $l$-th iteration if and only if at the $(l-1)$-th iteration, they have the same color and the same information aggregation from neighbors in terms of multiset of colors and edge weights. This is a color refinement procedure. One can have a partition of the vertex set $V\cup W$ at each iteration based on vertices’ colors: two vertices are classified in the same class if and only if they are of the same color. Such a partition is strictly refined in the first $\mathcal{O}(m+n)$ iterations and will remain stable or unchanged afterward if no collision, see e.g. [chen2022representing-lp]*Appendix A. We provide an explicit example of the color refinement and the partition generated by WL test in Figure 6 of the Appendix.
Intuitively, vertices in the same class of the final stable partition generated by the WL test will always have identical features in message-passing layers for all GNNs in $\mathcal{F}_{\textup{MI-LCQP}}$ or $\mathcal{F}_{\textup{MI-LCQP}}^{W}$, and vice versa, since the color refinement procedure in Algorithm 1 follows the same mechanism as the message-passing process. Therefore, for identifying a subset of $\mathcal{F}_{\textup{MI-LCQP}}$ on which GNNs have sufficiently strong separation power, we propose the following definition, which basically states that vertices in the same class generated by the WL test can indeed be treated same in some sense.
Definition 4.4 (Message-passing-tractable MI-LCQP).
In the same setting as in Definition 4.8, we say that $G_{\text{MI-LCQP}}$ is message-passing-tractable (MP-tractable) if the followings hold:
  1. (a)
    For any $p\in\{1,2,\dots,s\}$ and $q\in\{1,2,\dots,t\}$, $A_{ij}$ is constant in $i\in I_{p},j\in J_{q}$.
  2. (b)
    For any $q,q^{\prime}\in\{1,2,\dots,t\}$, $Q_{jj^{\prime}}$ is constant in $j\in J_{q},j^{\prime}\in J_{q^{\prime}}$.
We use $\mathcal{G}_{\textup{MP}}^{m,n}\subset\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$ to denote the collection of all MP-tractable MI-LCQP-graphs with $m$ constraints and $n$ variables.
The analysis in [chen2022representing-lp]*Appendix A suggests that all rows/columns of the submatrix $(A_{ij})_{i\in I_{p},j\in J_{q}}$ (or $(Q_{jj^{\prime}})_{j\in J_{q},j^{\prime}\in J_{q^{\prime}}}$) are equal as multisets of entries, and the MP-traceability further requires that all elements in those multisets are the same. Definition 4.4 is generalized from prior work on MILPs [chen2024rethinking] and with the assumption of MP-tractability, the universal approximation results can be established for GNNs to represent $\Phi_{\textup{feas}}$ and $\Phi_{\textup{obj}}$.
@@
$\mathbb{P}$ is a Borel regular probability measure on $\mathcal{G}^{m,n}_{\textup{MI-LCQP}}$  [^3] The topology of $\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$ is defined in the same way as $\mathcal{G}_{\textup{LCQP}}^{m,n}$..
@@
Let $\mathbb{P}$ be a probability measure on $\mathcal{G}^{m,n}_{\textup{MI-LCQP}}$ satisfying Assumption 4.5 and $\mathbb{P}[G_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{MP}}^{m,n}]=1$. For any $\epsilon>0$, there exists $F\in\mathcal{F}_{\textup{MI-LCQP}}$ such that
$\mathbb{P}\big{[}\mathbb{I}_{F(G_{\textup{MI-LCQP}})>\frac{1}{2}}\neq\Phi_{\textup{feas}}(G_{\textup{MI-LCQP}})\big{]}<\epsilon.$
@@
Let $\mathbb{P}$ be a probability measure on $\mathcal{G}^{m,n}_{\textup{MI-LCQP}}$ satisfying Assumption 4.5 and $\mathbb{P}[G_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{MP}}^{m,n}]=1$. The followings are true:
  1. (1)
    For any $\epsilon>0$, there exists $F_{1}\in\mathcal{F}_{\textup{MI-LCQP}}$ such that
    $\mathbb{P}\big{[}\mathbb{I}_{F_{1}(G_{\textup{MI-LCQP}})>\frac{1}{2}}\neq\mathbb{I}_{\Phi_{\textup{obj}}(G_{\textup{MI-LCQP}})\in\mathbb{R}}\big{]}<\epsilon.$
  2. (2)
    If $\mathbb{P}[\Phi_{\textup{obj}}(G_{\textup{MI-LCQP}})\in\mathbb{R}]=1$, then for any $\epsilon,\delta>0$, there exists $F_{2}\in\mathcal{F}_{\textup{MI-LCQP}}$ such that
    $\mathbb{P}\left[|F_{2}(G_{\textup{MI-LCQP}})-\Phi_{\textup{obj}}(G_{\textup{MI-LCQP}})|>\delta\right]<\epsilon.$
We then discuss the expressive power of GNNs to represent optimal solutions to MI-LCQP problems. Different from the LCQP setting, the optimal solution to an MI-LCQP problem may not exist even if it is feasible and bounded, i.e., $\Phi_{\textup{obj}}(G_{\textup{MI-LCQP}})\in\mathbb{R}$. Thus, we have to work with $\mathcal{G}_{\textup{sol}}^{m,n}\subset\Phi_{\textup{obj}}^{-1}(\mathbb{R})\subset\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$ where $\mathcal{G}_{\textup{sol}}^{m,n}$ is the collection of all MI-LCQP-graphs for which an optimal solution exists. For $G_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{sol}}^{m,n}$, it is possible that it admits multiple optimal solution. Moreover, there may even exist multiple optimal solutions with the smallest $\ell_{2}$-norm due to its non-convexity, which means that we cannot define the optimal solution mapping $\Phi_{\textup{sol}}$ using the same approach as in the LCQP case. To uniquely define $\Phi_{\textup{sol}}$, we need the following concept.
Definition 4.8 (Unfoldable MI-LCQP).
Let $G_{\text{MI-LCQP}}\in\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$ be a MI-LCQP problem and let $(\mathcal{I},\mathcal{J})$ be the final stable partition of $V\cup W$ generated by WL test without collision, where =$\mathcal{I}=\{I_{1},I_{2},\dots,I_{s}\}$ is a partition of $V=\{1,2,\dots,m\}$ and $\mathcal{J}=\{J_{1},J_{2},\dots,J_{t}\}$ is a partition of $W=\{1,2,\dots,n\}$. We say that $G_{\text{MI-LCQP}}$ is unfoldable if $t=n$ and $|J_{1}|=|J_{2}|=\cdots=|J_{n}|=1$, i.e., all vertices in $W$ have different colors. We use $\mathcal{G}_{\textup{unfold}}^{m,n}\subset\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$ to denote the collection of all unfoldable MI-LCQP-graphs with $m$ constraints and $n$ variables.
The MI-LCQP unfoldability is a generalization of the MILP unfoldability proposed in [chen2022representing-milp]. If we assume that $G_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{sol}}^{m,n}$ is unfoldable, then using the same approach as in [chen2022representing-milp]*Appendix C, one can define a total ordering on the optimal solution set and hence define $\Phi_{\textup{sol}}(G_{\textup{MI-LCQP}})$ as the minimal element in the optimal solution set, which is unique and permutation-equivariant, meaning that if one relabels vertices of $G_{\textup{MI-LCQP}}$, then entries of $\Phi_{\textup{sol}}(G_{\textup{MI-LCQP}})$ are relabelled accordingly. It can be proved that the unfoldability implies the MP-tractability but there exist MP-tractable instances that are not unfoldable (see Appendix C).
With the assumption that the MI-LCQP data distribution is supported in $\mathcal{G}_{\textup{sol}}^{m,n}\cap\mathcal{G}_{\textup{unfold}}^{m,n}$, a positive result for GNNs to represent the optimal solution mapping can be derived.
Theorem 4.9.
Let $\mathbb{P}$ be a probability measure on $\mathcal{G}^{m,n}_{\textup{MI-LCQP}}$ satisfying Assumption 4.5 and $\mathbb{P}[G_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{sol}}^{m,n}\cap\mathcal{G}_{\textup{unfold}}^{m,n}]=1$. For any $\epsilon,\delta>0$, there exists $F_{W}\in\mathcal{F}_{\textup{MI-LCQP}}^{W}$ such that
@@
The proofs for Theorems 4.6, 4.7, and 4.9 can be found in Appendix D. As the concluding remarks for this section, we would like to mention that although there exist MI-LCQP problems that are not MP-tractable/unfoldable, the MP-tractable/unfoldable MI-LCQP class is already rich enough to cover many applications. In particular, it can be proved MI-LCQP-graphs are generically MP-tractable/unfoldable (see Section C for details).

### v2

@@ 4.1. GNNs cannot universally represent MI-LCQPs
In this subsection, we answer the question (1.1) for MI-LCQP. When integer variables are introduced, the situation changes. Particularly, we present some counter-examples illustrating the fundamental limitation of GNNs in this context.
Proposition 4.1.
@@
It is indicated that for some MI-LCQP data distribution, it is impossible to train a GNN to predict MI-LCQP properties, regardless of the GNN’s size. Particularly, one can choose the uniform distribution over pairs of instances satisfying Propositions 4.1, 4.2, and 4.3: any GNN making good approximation on one instance must fail on the other.
The detailed proofs are provided in Appendix B. Here we present a pair of MI-LCQPs that prove Proposition 4.3.
A counterexample
We modify the two LCQPs in (3.3) and (3.4) into two MI-LCQPs by introducing integer constraints on all variables: specifically, $x_{j}\in\mathbb{Z}$ for $i\in\{1,\cdots,7\}$. Consequently, the node attributes $w_{j}$ are updated to $w_{j}=(1,0,3,1)$, where the last entry, $1$, indicates the integral constraint $\delta_{I}(j)=1$. All other components of the graphs remain unchanged, as described in Section 3. With this modification, the node attributes $w_{j}$ remain identical for all $j\in W$. Since all other graph components are unchanged, the same argument from Section 3 applies: any GNNs cannot distinguish the two MI-LCQPs.
However, the mixed-integer setting differs from the continuous one: the two problems no longer share the same solution $(\frac{6}{7},\frac{6}{7},\frac{6}{7},\frac{6}{7},\frac{6}{7},\frac{6}{7},\frac{6}{7})$. Instead, they have disjoint optimal solution sets. The first instance has an unique feasible (and thus optimal) solution (3,3,0,0,0,0,0){(3,3,0,0,0,0,0)}, while in the second instance, it is (2,2,2,0,0,0,0){(2,2,2,0,0,0,0)}.
@@ 4.2. GNNs can represent particular types of MI-LCQPs
We have shown a fundamental limitation of GNNs to represent MI-LCQP in general. A natural question arise: Whether we can identify a subset of $\mathcal{G}_{\textup{MI-LCQP}}$ on which it is possible to train reliable GNNs. To address this, we need to gain a better understanding for the separation power of GNNs, or equivalently, of the WL test, according to the discussion following Theorem 3.4. We state in Algorithm 1 the WL test for MI-LCQP-graphs associated to $\mathcal{F}_{\textup{MI-LCQP}}$ or $\mathcal{F}_{\textup{MI-LCQP}}^{W}$.
Algorithm 1 The WL test (Ref. to App. D for an example)
[…]
2:Initialization: $C_{i}^{0,V}=\text{hash}(v_{i})$, and $C_{j}^{0,W}=\text{hash}(w_{j})$.
@@
4:  $C_{i}^{l,V}=\text{hash}\left(C_{i}^{l-1,V},\{\!\!\{(C_{j}^{l-1,W},A_{ij})\}\!\!\}_{j\in\mathcal{N}_{i}^{W}}\right)$.
5:  $C_{j}^{l,W}=\text{hash}\left(C_{j}^{l-1,W},\{\!\!\{(C_{i}^{l-1,V},A_{ij})\}\!\!\}_{i\in\mathcal{N}_{j}^{V}},\{\!\!\{(C_{j^{\prime}}^{l-1,W},Q_{jj^{\prime}})\}\!\!\}_{j^{\prime}\in\mathcal{N}_{j}^{W}}\right).$
@@
7:return All vertex colors $\{\!\!\{C_{i}^{L,V}\}\!\!\}_{i=0}^{m},\{\!\!\{C_{j}^{L,W}\}\!\!\}_{j=0}^{n}$.
Here, $C_{i}^{l,V}$ and $C_{j}^{l,W}$ are the colors of node $i\in V$ and node $j\in W$ at the $l$-th iteration. The “hash” function is any injective (collision-free) mapping, independent of $i,j$. In practice, standard built-in hash functions can be used.
The WL test mimics GNNs’ iterative updates, replacing learnable mappings with hash functions. While hash functions cannot directly map a graph to desired outcomes, they effectively distinguish nodes and differentiate graphs.
@@
Initially, each vertex is labeled a color by a hash function according to its attributes ($v_{i}$ or $w_{j}$). At the $l$-th iteration, two vertices are of the same color if and only if at the (l1)(l-1)-th iteration, they have the same color and the same information aggregation from neighbors. This is a color refinement procedure. One can have a partition of the vertex set $V\cup W$ at each iteration based on vertices’ colors: two vertices are classified in the same class if and only if they are of the same color. Such a partition is strictly refined in the first $\mathcal{O}(m+n)$ iterations and will remain stable or unchanged afterward, see e.g., [berkholz2017tight].
Revisit the counterexample in Section 4.1
As analyzed earlier, applying the WL test to the two MI-LCQPs results in all variable nodes $W$ having identical colors (red nodes), meaning they belong to the same class in the WL test’s final stable partition. Nodes in the same class of this partition will always share identical attributes across all layers of any GNNs, and vice versa, because the color refinement process in Algorithm 1 mirrors GNNs’ update mechanisms. However, while the desired GNN output is $(3,3,0,0,0,0,0)$, GNNs cannot differentiate between these variables, resulting in outputs where $y_{1}=\cdots=y_{7}$. This contradiction is the core reason for GNNs’ failure on these counterexamples.
To address this issue, we propose to: (I) Restrict the focus to MI-LCQPs where each node in the graph representation has a unique color (i.e., every class in the partition contains exactly one node), ensuring GNNs can separate all nodes. or (II) Allow partitions with classes containing multiple nodes but impose additional conditions within each class to ensure the desired output does not rely on distinguishing between these nodes. Therefore, we define as follows.
Definition 4.4 (GNN-friendly MI-LCQPs).
For a MI-LCQP problem $G_{\text{MI-LCQP}}\in\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$, apply the WL test to obtain the final stable partition $(\mathcal{I},\mathcal{J})$, where $\mathcal{I}=\{I_{1},I_{2},\dots,I_{s}\}$ partitions $V=\{1,2,\dots,m\}$ and $\mathcal{J}=\{J_{1},J_{2},\dots,J_{t}\}$ partitions $W=\{1,2,\dots,n\}$. Then,
  • $G_{\text{MI-LCQP}}$ is called “GNN-solvable” if $t=n$ and $|J_{1}|=|J_{2}|=\cdots=|J_{n}|=1$, i.e., all vertices in $W$ have different colors. We use $\mathcal{G}_{\textup{solvable}}^{m,n}\subset\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$ to denote the collection of all GNN-solvable MI-LCQPs.
  • $G_{\text{MI-LCQP}}$ is called “GNN-analyzable” if: (1) For any $p\in\{1,2,\dots,s\}$ and $q\in\{1,2,\dots,t\}$, $A_{ij}$ is constant across all $i\in I_{p},j\in J_{q}$. (2) For any $q,q^{\prime}\in\{1,2,\dots,t\}$, $Q_{jj^{\prime}}$ is constant across $j\in J_{q},j^{\prime}\in J_{q^{\prime}}$. The set of all such MI-LCQPs is denoted by $\mathcal{G}_{\textup{analyzable}}^{m,n}\subset\mathcal{G}_{\textup{MI-LCQP}}^{m,n}$.
GNN-solvability extends “unfoldability” in the context of MILP [chen2022representing-milp], and GNN-analyzability extends “MP-tractability” [chen2024rethinking]. With such definitions, we can establish GNNs’ expressive power for MI-LCQPs.
@@
$\mathbb{P}$ is a Borel regular probability measure defined on the space of MI-LCQP-graphs $\mathcal{G}^{m,n}_{\textup{MI-LCQP}}$.
@@
For any $\mathbb{P}$ satisfying Assumption 4.5 and $\mathbb{P}[G_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{analyzable}}^{m,n}]=1$, and for any $\epsilon>0$, there exists $F\in\mathcal{F}_{\textup{MI-LCQP}}$ such that
(4.1) $\mathbb{P}\big[\mathbb{I}_{F(G_{\textup{MI-LCQP}})>\frac{1}{2}}\neq\Phi_{\textup{feas}}(G_{\textup{MI-LCQP}})\big]<\epsilon,$
and there exists $F_{1}\in\mathcal{F}_{\textup{MI-LCQP}}$ such that
(4.2) $\mathbb{P}\big[\mathbb{I}_{F_{1}(G_{\textup{MI-LCQP}})>\frac{1}{2}}\neq\mathbb{I}_{\Phi_{\textup{obj}}(G_{\textup{MI-LCQP}})\in\mathbb{R}}\big]<\epsilon.$
Additionally, if $\mathbb{P}[\Phi_{\textup{obj}}(G_{\textup{MI-LCQP}})\in\mathbb{R}]=1$, for any $\epsilon,\delta>0$, there exists $F_{2}\in\mathcal{F}_{\textup{MI-LCQP}}$ such that
(4.3) $\mathbb{P}\left[|F_{2}(G_{\textup{MI-LCQP}})-\Phi_{\textup{obj}}(G_{\textup{MI-LCQP}})|>\delta\right]<\epsilon.$
To extend these results to predicting optimal solutions, we need to assume the MI-LCQPs have an optimal solution. We denote $\mathcal{G}_{\textup{sol}}^{m,n}$ as the set of MI-LCQPs with an optimal solution. The assumption is expressed as $G_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{sol}}^{m,n}$.
@@
For any $\mathbb{P}$ satisfying Assumption 4.5 and $\mathbb{P}[G_{\textup{MI-LCQP}}\in\mathcal{G}_{\textup{sol}}^{m,n}\cap\mathcal{G}_{\textup{solvable}}^{m,n}]=1$, and for any $\epsilon,\delta>0$, there exists $F_{W}\in\mathcal{F}_{\textup{MI-LCQP}}^{W}$ such that
@@
Theorems 4.6 and 4.7 indicate the subsets of MI-LCQPs where GNNs can succeed: For GNN-analyzable MI-LCQPs, GNNs can approximate their feasibility and optimal objective; For GNN-solvable ones, GNNs can approximate their optimal solutions. Proofs are available in Appendix C.
4.3. Discussions of “friendly” MI-LCQPs
To better illustrate the practical implications of Theorems 4.6 and 4.7, we make some discussions of GNN-analyzability and GNN-solvability here.
Analyzability vs Solvability
While all GNN-solvable MI-LCQPs must be GNN-analyzable, not all GNN-analyzable problems are necessarily GNN-solvable. Related discussions, proofs and examples are detailed in Appendix D.
Since solvability is a stronger condition than analyzability, when a problem is GNN-solvable, GNNs can approximate not only the optimal solution (Theorem 4.7) but also the feasibility and the optimal objective (Theorem 4.6). In this sense, learning the solution mapping is a stricter requirement than learning the feasibility and optimal objective.
Frequency of friendly instances
In practice, the frequency of GNN-analyzable and GNN-solvable instances largely depends on the the level of symmetry in the dataset. Here, two variables labeled with the same color by the WL test are said to be symmetric. When there is symmetry in a MI-LCQP, it becomes GNN-insolvable; and higher symmetry increases the risk of being GNN-inanalyzable. Examples provided by (3.3) and (3.4) admit strong symmetry across all variables, making them neither GNN-analyzable nor GNN-solvable. Fortunately, GNN-solvable and GNN-analyzable instances make up the majority of the MI-LCQP set, under specific distributional assumptions (see e.g. Proposition D.3). This explains the observed success of GNNs for QPs in the existing literature and addresses the gap between theoretical limitations and empirical success discussed in Section 1. The dataset used in Section 5 consists entirely of GNN-solvable and GNN-analyzable instances. However, in some challenging, artificially collected or created datasets, such as QPlib [furini2019qplib], it’s possible to be GNN-insolvable or even GNN-inanalyzable (see in Appendix D.2 for details).
Numerical verification
While GNN-analyzability and GNN-solvability depend on the dataset, they can be efficiently verified in practice. One may first apply Algorithm 1, which requires at most $\mathcal{O}(m+n)$ iterations. The single-iteration complexity is bounded by the number of edges in the graph [shervashidze2011weisfeiler], which, in our context, is the number of nonzeros in matrices $A$ and $Q$: $\text{nnz}(A)+\text{nnz}(Q)$. Therefore, Algorithm 1’s complexity is $\mathcal{O}((m+n)\cdot(\text{nnz}(A)+\text{nnz}(Q)))$. Afterward, both conditions can be directly verified using Definition 4.4.
Therefore, these criteria provide practitioners with a practical tool to evaluate datasets before applying GNNs. Additionally, if challenges arise during GNN training on MI-LCQPs, verifying them can help identify potential issues.